40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Dec 17 14:02:05 2015, gautam, Update, CDS, IPC channels for beat frequency control set up 
    Reply  Thu Dec 17 16:44:03 2015, gautam, Update, CDS, ALS Slow control MEDM screen updated ALS_SLOW_17DEC2015.png
Message ID: 11891     Entry time: Thu Dec 17 16:44:03 2015     In reply to: 11890
Author: gautam 
Type: Update 
Category: CDS 
Subject: ALS Slow control MEDM screen updated 
Quote:

I've not updated the MEDM screens to reflect the two new paths yet, but will do so soon. It also remains to install appropriate filters for the servo path that takes the frequency readout as the input.

A few more related changes:

  1. The couplers that used to sit on the green beat PDs on the PSL table have now been shifted to the IR broadband PDs in the FOL box so that I can get the IR beat frequency over to the frequency counters. The FOL box itself, along with the fibers that bring IR light to the PSL table, have been relocated to the corner of the PSL table where the green beat PDs sit because of cable length constraints.
  2. I've updated the ALS slow control MEDM screen to allow for slow control of the beat frequency. The servo shape for now is essentially just an integrator with a zero at 1 Hz. The idea is to set an offset in the new filter module, which is the desired beat frequency, and let the integrator maintain this beat frequency. One thing I've not taken care of yet is automatically turning this loop off when the IMC loses lock. Screenshot of the modified MEDM screen is attached. 
  3. I checked the performance by using the temperature sliders to introduce an offset. The integrator is able to bring the beat frequency back to the setpoint in a few seconds, provided the step I introduced was not two big (~20 counts, but this is a pretty large shift in beat frequency, nearly 20MHz).

To do:

  1. Figure out how to deal with the IMC losing lock. I guess this is important if we want to use the IR beatnote as a diagnostic for the state of the X AUX laser.
  2. Optimize the servo gains a little - I still see some ringing when I introduce an offset, this could be avoided...
Attachment 1: ALS_SLOW_17DEC2015.png  32 kB  | Hide | Hide all
ALS_SLOW_17DEC2015.png
ELOG V3.1.3-